An fm radio station broadcast at a frequency of 109.8 hz. What is the wavelength of the radio signal?

1
Expert's answer
2022-06-07T17:37:14-0400
"\lambda=\frac{c}{f}=\frac{3*10^8}{109.8}=2.73*10^6\:\rm m"